# How do language models learn facts? Dynamics, curricula and hallucinations

### v1

@@ Abstract
Large language models offer a powerful and intuitive interface to access the vast amounts of knowledge on which they were trained. The learning process acts as a lossy compression algorithm turning training data into neural network parameters, thereby implicitly determining what information is preserved within the final model. The extent and nature of this information loss depend on numerous factors, including architecture, training objective, and data distribution, dependencies which remain poorly understood. Deciphering the mechanisms underlying knowledge compression is increasingly important as these models are becoming our main gateway to human knowledge. Our work addresses this challenge by systematically investigating how language models acquire factual knowledge.
@@ 1 An experimental setup to track knowledge over the course of learning
This study investigates the learning dynamics underlying factual recall and knowledge acquisition in language models. This presents two main methodological challenges: First, we need to isolate the model’s knowledge from other abilities. Second, we want to evaluate the models over the course of learning, which requires computationally efficient measurement techniques. Following Allen-Zhu and Li (2023), we train language models on synthetic biographies that feature key properties of datasets used to train large language models. By carefully designing these synthetic biographies, we can attribute the model’s ability to predict specific tokens solely to its acquired knowledge and efficiently measure its knowledge through its performance on these tokens. In this section, we first define knowledge and contrast it with memory, then describe our synthetic dataset and introduce the metrics we use to track knowledge during training, and finally describe training details.
@@ 1.2 Synthetic biographies as a framework for studying knowledge
We summarize the data generation process in Figure 1. The dataset contains a population of $N$ randomly sampled individuals, each with a unique name and six attributes: birthdate, birthplace, university, major, company, and location. To generate a biography, we first sample an individual from the population. For each of attribute, we then sample a template from a finite pool, fill it with the individual’s information, and concatenate the resulting sentences in random order to form a complete biography. It should be noted that to make the dataset suitable for our study, our template generation and manipulation differs from the original setup. We detail these modifications in the next section and provide additional details in the appendix.
@@ 1.3 Measuring knowledge at scale
Measuring knowledge in language models typically relies on factual question-answering tasks (Lin et al., 2021). However, this approach is computationally intractable in our setup, as it requires repeatedly fine-tuning models to handle question-answer prompt formats throughout the training process. Our framework offers a simple solution to this challenge: by consistently placing information about the individual and attribute type before the attribute value, we transform each attribute value prediction into a factual recall task. This lets us quantify knowledge via the model’s loss and accuracy on predicting attribute value tokens. We denote them as attribute loss and attribute accuracy. Formal definitions are provided in the appendix. The attribute accuracy metric is arguably more intuitive, as it serves as a proxy for the accuracy at which the model, once fine-tuned, would correctly answer recall questions. However, we focus most of our analysis on the attribute loss due to its greater interpretive power and as it generally exhibits more consistent progression across various scales (Schaeffer et al., 2024).
Two important considerations remain to be solved: how to distinguish knowledge from memorization and what is a meaningful baseline for the attribute loss. To address the first point, we evaluate models on unseen biographies of previously encountered individuals. We do so by randomly splitting, for each individual, the template pool into training and evaluation templates, with this split varying across individuals. This ensures that models are evaluated on entirely new biography formulations, thereby ruling out the possibility that strong performance stems from memorization. For the second point, we introduce a theoretical baseline corresponding to the best possible performance achievable by a model with no individual-specific knowledge. Such a model would only know the overall distribution of attribute values, and its loss would equal the entropy of the attribute value distribution. We refer to this value as the no knowledge baseline. Any model performing better than this baseline must necessarily have acquired some knowledge about specific individuals in the training distribution.
@@ 1.4 Language models are trained following standard recipes
While we use synthetic data in our experiments, we keep the architecture and the training protocol as close as possible to standard practices. By default, we train an 8-layer decoder-only Transformer ($44$M non-embedding parameters) with the same structure as in Hoffmann et al. (2022) using the AdamW optimizer (Loshchilov and Hutter, 2019) and a cosine learning rate schedule without warm-up. We tune the learning rate in all our experiments. Additional details are provided in the appendix.
@@ 2 How language models acquire knowledge during learning
Figure 2: Knowledge acquisition occurs in three phases. (left) In a very short first phase, the model learns generic attribute value statistics. In the second phase, performance plateaus at a level achievable by an ideal model lacking individual-specific knowledge (this corresponds to the no knowledge baseline and a near-zero knowledge accuracy). This plateau’s duration is nearly proportional to the number of individuals (right). Finally, the model learns associations between subjects and attributes: knowledge emerges as the model is trained longer (middle). Results are averaged over 5 seeds ($\pm$ std). See Section 2.1 for details.
@@ 2.1 The three phases underlying knowledge acquisition
The temporal evolution of attribute loss (Figure 2) exhibits a consistent three-phase pattern:
@@
    Knowledge emergence. The model finally enters a knowledge acquisition phase, in which it progressively learns association between individuals and their attributes. Its knowledge accuracy progressively leaves the near-zero regime in this phase: knowledge emerges.
The pattern described above is robust across a range of hyperparameter configurations. Specifically, qualitative results remain consistent when varying learning rate, weight decay, batch size, number of individuals, model size, and even the type of sequence mixing block (attention-based vs. recurrent), cf. Figure C. This is consistent with the findings of Nichani et al. (2024), who observed a similar pattern in a simplified, theoretically-tractable version of the task learned by a linear attention layer. These findings suggest that the data structure, particularly the high individual-to-attribute ratio, plays a critical role in driving this multi-phase learning behavior. We additionally discuss connections to existing work studying Transformer learning dynamics through the lens of $N$-grams in Appendix A.
@@ 2.2 The attention-based circuits supporting recall are created during the loss plateau
Previous work has analyzed how Transformer-based language models solve factual recall tasks similar to the one we are interested in this study. It can be summarized as follows (cf. Figure E for a visual summary). The first attention layer aggregates name tokens together to form a representation of the individual’s name at the position of the last name token. This representation serves as a query for subsequent layers, particularly the multi-layer perceptrons, which effectively act as a key-value database (Geva et al., 2021; Meng et al., 2022). Individual-specific knowledge thus typically appears within the final name token’s residual stream (Nanda et al., 2023b; Allen-Zhu and Li, 2023; Geva et al., 2023). The final attention layer then selects the relevant information conditioned on the attribute type and makes it available for predicting attribute value tokens when needed (Geva et al., 2023; Nanda et al., 2023b). With this last circuit established, attribute value prediction errors directly backpropagate to relevant tokens, here the name tokens. However, without it, errors are spread across irrelevant tokens, hindering learning efficiency. We therefore hypothesize that the plateau phase corresponds to the formation of these attention-based recall circuits, with learning speed directly linked to their development stage.
@@
Figure 3: The attention-based circuits supporting recall are created during the loss plateau. (left) We design an attention patching experiment, in which we take a snapshot of a reference model at some time during its training, and use its attention patterns as a replacement for the ones of a modified model throughout its training. (middle) The more trained the reference model is, the better its attention patterns are for the modified model and these changes mainly happen during the plateau. The very beginning of learning is an exception to this trend. This correlates with the fact that, at this time during training, the name tokens (compared to the rest of the text, which contains information about the attribute type) receive reduced attention when the first attribute value token is predicted (cf. right panel). See Section 2.2 for more detail.
@@
A consequence of that hypothesis is that a model with learned attention patterns should learn faster than one with pre-learning patterns. To test this, we design the following attention patching experiment (Figure 3, left). We first train a reference model, and save reference checkpoints from it at various stages of training. We then restart training from scratch, but replace the model’s attention patterns [^1] We call attention patterns the softmax of the attention logits. with those produced from one of the reference checkpoints. We expect the quality of the attention patterns, measured by how easy it is to learn the task with them, provided by the reference model to progressively improve as they are taken closer to the plateau end. This is what we observe empirically (Figure 3, middle), and the plateau disappears when providing the model with learned attention patterns (i.e. post-plateau patterns). Interestingly, the attention patterns acquired early during learning (e.g., at steps $500$ and $1$k) are significantly worse than the pre-learning ones, likely due to initial focus on predicting the attribute value distribution conditioned on the attribute type only. This attention on attribute type tokens rather than on name tokens slows down learning, as per our argument from the last paragraph.
@@
So far, our results show the formation of an attention-based circuit during the plateau phase, though not necessarily the specific extraction circuit we hypothesized. To investigate this further, we examine the evolution of the network’s attention patterns during training for signatures of such a circuit. Specifically, we analyze the attention paid to name tokens (relative to general text tokens containing attribute type information) when predicting the first attribute value token, the primary position testing factual recall. We observe this attention increasing throughout the plateau, after being significantly lower during the initial phase where predictions align with the generic attribute value distribution (Figure 3, right panel). This offers an intuitive explanation for the qualitative differences in attention patterns observed in Figure 3 (middle). A similar analysis for the name tokens grouping circuit, characterized by high attention to name tokens from the last name token in the first layer, reveals its emergence within a few hundred steps. This may be specific to our setup, in which such a circuit is useful to better predict which template comes next. Details of this analysis are provided in Appendix D.2. Together with our attention patching results, this result suggests that the attention-based extraction circuit responsible for recall develops during the plateau phase.
@@ 3 Data distributional properties drive knowledge acquisition
Figure 4: Data distributional properties can speed up knowledge acquisition. (left) The length of the plateau significantly decreases when some individuals appear more frequently than other, which is here achieved by increasing $\alpha$. (middle) As a result, it is beneficial to train the model on more imbalanced distributions, particularly as the number of training steps decreases or as the total number of individuals increases. (right) Such a strategy, improves the final amount of knowledge contained in the network (purple vs. grey line). Dynamically adapting the data distribution yields even larger benefits (blue line). See Section 3 for more detail.
@@ 3.1 The trade-off underlying imbalances in the data distribution
Building on the analysis presented in the preceding section, we can already develop an intuition of how data distribution impact the plateau and knowledge acquisition phases. Assuming that the circuits the network creates during the plateau transfer to other individuals, reducing the number of individuals should minimize time spent in this phase, as shown in Figure 2 (right). Naturally this assumption will break when an excessively small number of individuals is over represented, as the model will likely overfit. As a consequence, imbalanced distributions with a small group of over represented individuals should generally shorten the plateau duration. However, this does not hold for the knowledge acquisition phase. Each individual contributes equally to the total amount of knowledge the model has, so the attribute loss on the entire population will asymptotically behave like the one on the group which is learned slowly, that is the least frequent one. For that reason, a uniform distribution should be optimal for knowledge acquisition. As a consequence, there is a trade-off: imbalanced distributions speed up the plateau and slow down knowledge acquisition whereas the reverse holds for more uniform distributions. The distribution that optimizes the final amount of knowledge stored within the network weights should therefore become increasingly more imbalanced as the plateau takes a larger portion of training time, i.e., when the number of training steps decreases or when the number of individuals increases.
@@
We are aware of three related findings. First, Allen-Zhu and Li (2023) demonstrated the benefits of “celebrities” in a similar setup to ours. Our analysis provides a partial explanation for this phenomenon. While we find that celebrities offer comparable benefits to our inverse power law distribution (cf. Appendix E), their effect may be amplified in their setup due to their use of a single biography for non-celebrities, whereas we never repeat biographies. Second, Charton and Kempe (2024) showed that repetition benefits Transformer training on arithmetic tasks. Their learning curves exhibit numerous plateaus, and sample-level repetition reduces time spent on these plateaus. Our analysis may generalize to this task, suggesting a broader principle underlying our findings. Finally, Park et al. (2024) trained Transformers on a synthetic mixture of Markov chains found that low task diversity shortens plateau length, albeit leading to solutions that do not generalize as well out-of-distribution.
@@ 3.2 Data schedulers increase the final amount of acquired knowledge
We implement a “warm-up” strategy: the model initially trains on a subset of individuals for a fixed number of steps, and on all individuals afterwards. Starting with a subset of individuals should shorten the plateau, while the subsequent uniform distribution maximizes knowledge acquisition once the recall circuits are well established. Indeed, we observe significant gains, particularly when the number of individuals is large and the model starts acquiring knowledge (Figure 4, right), which is a practically relevant regime. Experimental details and additional analysis are available in the appendix.
@@ 4 Hallucinations hinder the integration of new knowledge post-training
This final section examines the challenge of expanding language models’ parametric knowledge through post-training procedures like fine-tuning. We find that certain types of hallucinations (overconfident predictions on unseen individuals) emerge simultaneously with knowledge about individuals within the training distribution and can be detected at the population level. These hallucinations significantly impact learning dynamics on new data, requiring numerous training steps to overcome miscalibration, during which pre-existing knowledge is significantly degraded. Adding replay of existing knowledge to the fine-tuning data mix only partially mitigates this issue. Overall, our findings offer an explanation for the infrequent use of fine-tuning (Ovadia et al., 2023; Jain et al., 2024) for incorporating new knowledge in the model’s parameters.

### v2

@@ Abstract
Large language models offer a powerful and intuitive interface to access the vast amounts of knowledge on which they were trained. The learning process acts as a lossy compression algorithm turning training data into neural network parameters, thereby implicitly determining what information is preserved within the final model. The extent and nature of this information loss depend on numerous factors, including architecture, training objective, and data distribution, dependencies that remain poorly understood. Deciphering the mechanisms underlying knowledge compression is increasingly important as these models are becoming our main gateway to human knowledge. Our work addresses this challenge by systematically investigating how language models acquire factual knowledge.
@@ 1 An experimental setup to track knowledge over the course of learning
This study investigates the learning dynamics underlying factual recall and knowledge acquisition in language models. This presents two main methodological challenges: First, we need to isolate the model’s knowledge from other abilities. Second, we want to evaluate the models over the course of learning, which requires computationally efficient measurement techniques. Following Allen-Zhu and Li (2023), we train language models on synthetic biographies that feature key properties of datasets used to train large language models. By carefully designing these synthetic biographies, we can attribute the model’s ability to predict specific tokens solely to its acquired knowledge and efficiently measure its knowledge through its performance on these tokens. In this section, we first define knowledge and contrast it with memory, then describe our synthetic dataset and introduce the metrics we use to track knowledge during training, and finally, describe training details.
@@ 1.2 Synthetic biographies as a framework for studying knowledge
We summarize the data generation process in Figure 1. The dataset contains a population of $N$ randomly sampled individuals, each with a unique name and six attributes: birthdate, birthplace, university, major, company, and location. To generate a biography, we first sample an individual from the population. For each attribute, we then sample a template from a finite pool, fill it with the individual’s information, and concatenate the resulting sentences in random order to form a complete biography. It should be noted that to make the dataset suitable for our study, our template generation and manipulation differ from the original setup. We detail these modifications in the next section and provide additional details in Appendix B.
@@ 1.3 Measuring knowledge at scale
Measuring knowledge in language models typically relies on factual question-answering tasks (Lin et al., 2021). However, this approach is computationally intractable in our setup, as it requires repeatedly fine-tuning models to handle question-answer prompt formats throughout the training process. Our framework offers a simple solution to this challenge: by consistently placing information about the individual and attribute type before the attribute value, we transform each attribute value prediction into a factual recall task. This lets us quantify knowledge via the model’s loss and accuracy on predicting attribute value tokens. We denote them as attribute loss and attribute accuracy. Formal definitions are provided in Appendix B. The attribute accuracy metric is arguably more intuitive, as it serves as a proxy for the accuracy with which the model, once fine-tuned, would correctly answer recall questions. However, we focus most of our analysis on the attribute loss due to its greater interpretive power and as it generally exhibits more consistent progression across various scales (Schaeffer et al., 2024).
Two important considerations remain to be solved: how to distinguish knowledge from memorization, and what is a meaningful baseline for the attribute loss. To address the first point, we evaluate models on unseen biographies of previously encountered individuals. We do so by randomly splitting the template pool into training and evaluation templates for each individual, with this split varying across individuals. This ensures that models are evaluated on entirely new biography formulations, thereby ruling out the possibility that strong performance stems from memorization. For the second point, we introduce a theoretical baseline corresponding to the best possible performance achievable by a model with no individual-specific knowledge. Such a model would only know the overall distribution of attribute values, and its loss would be equal to the entropy of the attribute value distribution. We refer to this value as the no knowledge baseline. Any model performing better than this baseline must necessarily have acquired some knowledge about specific individuals in the training distribution.
@@ 1.4 Language models are trained following standard recipes
While we use synthetic data in our experiments, we keep the architecture and the training protocol as close as possible to standard practices. By default, we train an 8-layer decoder-only Transformer ($44$M non-embedding parameters) with the same structure as in Hoffmann et al. (2022) using the AdamW optimizer (Loshchilov and Hutter, 2019) and a cosine learning rate schedule without warm-up. We tune the learning rate in all our experiments. Additional details are provided in Appendix B.
@@ 2 How language models acquire knowledge during learning
Figure 2: Knowledge acquisition occurs in three phases. (left) In a very short first phase, the model learns generic attribute value statistics. In the second phase, performance plateaus at a level achievable by an ideal model lacking individual-specific knowledge (this corresponds to the no knowledge baseline defined in Section 1.3 and a near-zero knowledge accuracy). This plateau’s duration is nearly proportional to the number of individuals (right). Finally, the model learns associations between subjects and attributes: knowledge emerges as the model is trained longer (middle). Results are averaged over 5 seeds ($\pm$ std). See Section 2.1 for details.
@@ 2.1 The three phases underlying knowledge acquisition
The temporal evolution of the attribute loss (Figure 2) exhibits a consistent three-phase pattern:
@@
    Knowledge emergence. The model finally enters a knowledge acquisition phase, in which it progressively learns associations between individuals and their attributes. Its knowledge accuracy progressively leaves the near-zero regime in this phase: knowledge emerges.
The pattern described above is robust across a range of hyperparameter configurations. Specifically, qualitative results remain consistent when varying learning rate, weight decay, batch size, number of individuals, model size, and even the type of sequence mixing block (attention-based vs. recurrent), cf. Figure C. This is consistent with the findings of Nichani et al. (2024), who observed a similar pattern in a theoretically tractable version of the task learned by a linear attention layer. These findings suggest that the data structure, particularly the high individual-to-attribute ratio, plays a critical role in driving this multi-phase learning behavior. Gu et al. (2025) found a similar behavior in a setup closely related to ours, and that, below some critical model size, the model does not learn any individual-specific information at all despite extensive training. We additionally discuss connections to existing work studying Transformer learning dynamics through the lens of $N$-grams in Appendix A.
@@ 2.2 The attention-based circuits supporting recall are created during the loss plateau
Previous work has analyzed how Transformer-based language models solve factual recall tasks similar to the one we are interested in this study. It can be summarized as follows (cf. Figure E for a visual summary). The first attention layers aggregate name tokens together to form a representation of the individual’s name at the position of the last name token. This representation serves as a query for subsequent layers, particularly the multi-layer perceptrons, which effectively act as a key-value database (Geva et al., 2021; Meng et al., 2022). Individual-specific knowledge thus typically appears within the final name token’s residual stream (Nanda et al., 2023b; Allen-Zhu and Li, 2023; Geva et al., 2023). The final attention layer then selects the relevant information conditioned on the attribute type and makes it available for predicting attribute value tokens when needed (Geva et al., 2023; Nanda et al., 2023b). With this last circuit established, attribute value prediction errors directly backpropagate to relevant tokens, here the name tokens. However, without it, errors are spread across irrelevant tokens, hindering learning efficiency. We therefore hypothesize that the plateau phase corresponds to the formation of these attention-based recall circuits, with learning speed directly linked to their development stage.
@@
Figure 3: The attention-based circuits supporting recall are created during the loss plateau. (left) We design an attention patching experiment, in which we take a snapshot of a reference model at some time during its training, and use its attention patterns as a replacement for the ones of a modified model throughout its training. (middle) The more trained the reference model is, the better its attention patterns are for the modified model, and these changes mainly happen during the plateau. The very beginning of learning is an exception to this trend. This correlates with the fact that, at this time during training, the name tokens (compared to the rest of the text, which contains information about the attribute type) receive reduced attention when the first attribute value token is predicted (cf. right panel). See Section 2.2 for more details.
@@
A consequence of that hypothesis is that a model with learned attention patterns should learn faster than one with pre-learning patterns. To test this, we design the following attention patching experiment (Figure 3, left). We first train a reference model and save reference checkpoints from it at various stages of training. We then restart training a model from scratch, but replace the model’s attention patterns [^1] We call attention patterns the softmax of the attention logits. with those produced from one of the reference checkpoints. This means that all the token-to-token interactions are specified by the reference model and that the modified model only learns token-wise feedforward transformations. We expect the quality of the attention patterns, measured by how easy it is to learn the task with them, provided by the reference model to progressively improve as they are taken closer to the plateau end. This is what we observe empirically (Figure 3, middle), and the plateau disappears when providing the model with learned attention patterns (i.e. post-plateau patterns). Interestingly, the attention patterns acquired early during learning (e.g., at steps $500$ and $1$k) are significantly worse than the pre-learning ones, likely due to initial focus on predicting the attribute value distribution conditioned on the attribute type only. This attention on attribute type tokens rather than on name tokens slows down learning, as per our argument from the last paragraph.
@@
So far, our results show the formation of an attention-based circuit during the plateau phase, though not necessarily the specific extraction circuit we hypothesized. To investigate this further, we examine the evolution of the network’s attention patterns during training for signatures of such a circuit. Specifically, we analyze the attention paid to name tokens (relative to general text tokens containing attribute type information) when predicting the first attribute value token, the primary position testing factual recall. We observe this attention increasing throughout the plateau, after being significantly lower during the initial phase where predictions align with the generic attribute value distribution (Figure 3, right panel). This offers an intuitive explanation for the qualitative differences in attention patterns observed in Figure 3 (middle). A similar analysis for the name tokens grouping circuit, characterized by high attention to name tokens from the last name token in the first layer, reveals its emergence within a few hundred steps. This may be specific to our setup, in which such a circuit is useful to better predict which template comes next. Details of this analysis are provided in Appendix D.2. Together with our attention patching results, this result suggests that the attention-based extraction circuit responsible for recall develops during the plateau phase, consistent with the findings of Ou et al. (2025). Note that our results do not rule out that other qualitative changes may occur during this phase, and they do not provide any mechanistic insights regarding how models acquire individual-specific knowledge.
@@ 3 Data distributional properties drive knowledge acquisition
Figure 4: Data distributional properties can speed up knowledge acquisition. (left) The length of the plateau significantly decreases when some individuals appear more frequently (up to some extent) than other, which is here achieved by increasing $\alpha$. (middle) As a result, it is beneficial to train the model on more imbalanced distributions (higher $\alpha$ values), particularly as the number of training steps decreases or as the total number of individuals increases. This panel reports the $\alpha$ value that minimizes the final attribute loss for each number of steps and individuals. (right) Such a strategy, improves the final amount of knowledge contained in the network (purple vs. grey line). Dynamically adapting the data distribution yields even larger benefits (blue line). See Section 3 for more detail.
@@ 3.1 The trade-off underlying imbalances in the data distribution
Building on the analysis presented in the preceding section, we can already develop an intuition of how data distribution impact the plateau and knowledge acquisition phases. Assuming that the circuits that the network creates during the plateau transfer to other individuals, reducing the number of individuals should minimize time spent in this phase, as shown in Figure 2 (right). Naturally, this assumption will break when an excessively small number of individuals is over-represented, as the model will likely overfit. As a consequence, imbalanced distributions with a small group of over represented individuals should generally shorten the plateau duration. However, this does not hold for the knowledge acquisition phase. Each individual contributes equally to the total amount of knowledge the model has, so the attribute loss on the entire population will asymptotically behave like the one on the group which is learned slowly, that is, the least frequent one. For that reason, a uniform distribution should be optimal for knowledge acquisition. As a consequence, there is a trade-off: imbalanced distributions speed up the plateau and slow down knowledge acquisition, whereas the reverse holds for more uniform distributions. The distribution that optimizes the final amount of knowledge stored within the network weights should therefore become increasingly more imbalanced as the plateau takes a larger portion of training time, i.e., when the number of training steps decreases or when the number of individuals increases.
@@
We are aware of three related findings. First, Allen-Zhu and Li (2023) demonstrated the benefits of “celebrities” in a similar setup to ours. Our analysis provides a partial explanation for this phenomenon. While we find that celebrities offer comparable benefits to our inverse power law distribution (cf. Appendix E), their effect may be amplified in their setup due to their use of a single biography for non-celebrities, whereas we never repeat biographies. Second, Charton and Kempe (2024) showed that repetition benefits Transformer training on arithmetic tasks. Their learning curves exhibit numerous plateaus, and sample-level repetition reduces time spent on these plateaus. Our analysis may generalize to this task, suggesting a broader principle underlying our findings. Finally, Park et al. (2024) trained Transformers on a synthetic mixture of Markov chains and found that low task diversity shortens plateau length, albeit leading to solutions that do not generalize as well out-of-distribution.
@@ 3.2 Data schedulers increase the final amount of acquired knowledge
We implement a “warm-up” strategy: the model initially trains on a subset of individuals for a fixed number of steps, and on all individuals afterwards. Starting with a subset of individuals should shorten the plateau, while the subsequent uniform distribution maximizes knowledge acquisition once the recall circuits are well established. Indeed, we observe significant gains, particularly when the number of individuals is large and the model starts acquiring knowledge (Figure 4, right), which is a practically relevant regime. Experimental details and additional analysis are available in Appendix E.
@@ 4 Hallucinations hinder the integration of new knowledge post-training
This final section examines the challenge of expanding language models’ parametric knowledge through post-training procedures like fine-tuning. We find that certain types of hallucinations (overconfident predictions on unseen individuals) emerge simultaneously with knowledge about individuals within the training distribution and can be detected at the population level. These hallucinations significantly impact learning dynamics on new data, requiring numerous training steps to overcome miscalibration, during which pre-existing knowledge is significantly degraded, a phenomenon reminiscent of catastrophic forgetting McCloskey and Cohen (1989). Adding replay of existing knowledge to the fine-tuning data mix only partially mitigates this issue. Overall, our findings offer an explanation for the infrequent use of fine-tuning (Ovadia et al., 2023; Jain et al., 2024) for incorporating new knowledge in the model’s parameters.